OpenStudy (anonymous):

Factoring? I need help factoring quadratic equations with numbers: mn - 15 + 3m - 5n

OpenStudy (anonymous):

\[mn+3m-5n-15=m(n+3)-5(n+3)=?\]

OpenStudy (anonymous):

Lol, thanks.

OpenStudy (mathmale):

So: (n+3) is the common factor. Please take this factoring to completion by factoring out (n+3).

OpenStudy (anonymous):

(mn + 3) (5n + 3)

OpenStudy (anonymous):

\[\left( n+3 \right)\left( m-5 \right)\]

OpenStudy (anonymous):

Thanks :)